# Extrinsic contribution to bosonic thermal Hall transport

### v1

@@ III.4 The disorder-induced curvature
and we note that a necessary condition for this small-deviation expansion to be well grounded is that $({\mathsf{D}}_{\rm d})_{n,n}(p)\neq 0\;\forall(n,p)$, i.e. that the diagonal elements of $\mathsf{S}^{-1}\mathsf{W}\mathsf{S}$ be non-vanishing. To the order we consider in $\lambda_{\rm bos}/\lambda_{\rm dis}\ll 1$ (small deviation expansion), in Eq.(31) $\mathsf{\Omega}^{\mathsf{W}}$ can be evaluated equivalently at $p$ or $p^{\prime}$. The quantity Eq.(32) may be dubbed for convenience the “disorder-induced curvature”, although it does not a priori qualify as a curvature in the strict mathematical sense. { In the scalar fermionic case, with $\mathsf{W}(p)\propto\text{\textbb{1}}$, the two curvatures coincide: then one just has
@@
We notice that in general $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}\neq-\mathsf{\Omega}^{\mathsf{W}}_{p_{\nu}p_{\mu}}$, at odds with the usual Berry curvature. Also, in Eq.(32) the sum of the first and second terms is gauge-invariant, while the third term by itself is gauge-invariant.
Besides our rigorous formal proof of Eq.(27) (and of the energy current below) including semiclassical disorder corrections, equations Eqs. (31), (32) are the main new result of this work, displaying both qualitative and quantitative differences with the known theory for the scalar fermionic case.
@@ IV.1 Solving the kinetic equation
×$\quad\times\delta\Big{(}(\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p^{\prime})-(\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p)\Big{)},$ (36)
@@ IV.3 The “side-jump” accumulation term
Expanding the Moyal product, we find to the order $O(\hbar^{1})$ two new contributions in the energy current density: one which we argue does not have a semiclassical interpretation and will henceforth be neglected, and one which we identify as the side-jump current,
@@
×$\qquad\qquad\qquad\times(p^{\prime}-p)_{\nu}\,(\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}})_{n,n}(p)$
×$\times\delta\Big{(}(\underline{\mathsf{K}}_{\rm d})_{n,n}(p^{\prime})-(\underline{\mathsf{K}}_{\rm d})_{n,n}(p)\Big{)}\,(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p^{\prime}).$
@@
×$\qquad\times\partial_{T}f((\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p))\;\partial_{p_{\lambda}}(\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p).$
@@ IV.4 Total contribution of disorder to the thermal Hall conductivity
We note that because $\mathsf{\Omega}^{\mathsf{W}}$ is not antisymmetric a priori, ${J}_{x}^{\rm dis}\neq 0$ a priori, so that from Eqs. (37),(40) there may be a $O(\hbar^{1})$ contribution to the longitudinal thermal conductivity, which is mainly $O(\hbar^{0})$.
IV.5 Analogous derivation for the electronic Hall conductivity
The same steps can be adapted with very little change to derive the extrinsic electrical Hall effect of electrons. Importantly, the normalization of eigenvectors is now $\mathsf{S}^{\dagger}\star\mathsf{S}=\text{\textbb{1}}$, so that $\underline{\mathsf{F}}_{\rm d}$ has dimension of a number not an energy. The equilibrium distribution is then $f(\varepsilon)=n_{\rm F}(\varepsilon-{\mathsf{e}}\phi)$, with $n_{\rm F}(x)=1/(e^{x/T}+1)$ the Fermi function, $-\mathsf{e}$ the electron charge and $\phi$ the electrical potential ($\bm{E}=-\bm{\nabla}\phi$ the electric field).
@@ IV.5 Analogous derivation for the electronic Hall conductivity
$\partial_{t}{\rm Tr}\,\mathsf{F}=\frac{1}{\hbar}{\rm Im}{\rm Tr}\left[\mathsf{K}+\mathsf{V}\;\overset{\star}{,}\;\mathsf{F}\right],$
@@
×$\times\left[{\mathsf{\Omega}}^{\mathsf{W}}_{p_{y}p_{\lambda}}(p)\,\partial_{p_{x}}\underline{\mathsf{K}}_{\rm d}(p)-{\mathsf{\Omega}}^{\mathsf{W}}_{p_{x}p_{\lambda}}(p)\,\partial_{p_{y}}\underline{\mathsf{K}}_{\rm d}(p)\right],$ (42)
see details in Appendix F.2. This formula generalizes the known result for electron side-jump based on a wavepacket analysis Sinitsyn et al. (2005), by allowing for impurities with a local matrix structure $\hat{W}$. In the “scalar fermionic” limit (where $\mathsf{\Omega}^{\mathsf{W}}=\mathsf{\Omega}$ reduces to the Berry curvature), assuming a quadratic dispersion and a uniform (i.e. momentum-independent) Berry curvature, one recovers immediately the familiar result $\sigma_{xy}^{\rm dis}=-2\sigma_{xy}^{\rm clean}$, where $\sigma_{xy}^{\rm clean}=-{\mathsf{e}}^{2}p_{\rm F}^{2}/4\pi$ is given by the Karplus-Luttinger formula ($p_{\rm F}$ the Fermi momentum).
@@ V Applications
We now consider an application to three slightly different physical models with disorder, the results of which are displayed in Fig. 1.
V.1 General bosonic model: physical picture

### v2

@@ III.4 The disorder-induced curvature
and we note that a necessary condition for this small-deviation expansion to be well grounded is that $({\mathsf{D}}_{\rm d})_{n,n}(p)\neq 0\;\forall(n,p)$, i.e. that the diagonal elements of $\mathsf{S}^{-1}\mathsf{W}\mathsf{S}$ be non-vanishing. To the order we consider in $\lambda_{\rm bos}/\lambda_{\rm dis}\ll 1$ (small deviation expansion), in Eq.(31) $\mathsf{\Omega}^{\mathsf{W}}$ can be evaluated equivalently at $p$ or $p^{\prime}$. Also, in Eq.(32) the sum of the first and second terms is gauge-invariant, while the third term by itself is gauge-invariant. An explicitly gauge-invariant formula for $\tfrac{1}{2}(\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}-\mathsf{\Omega}^{\mathsf{W}}_{p_{\nu}p_{\mu}})$, suitable for numerical evaluation, is provided in App. G.
The quantity Eq.(32) may be dubbed for convenience the “disorder-induced curvature”, although it does not a priori qualify as a curvature in the strict mathematical sense. { In the case of fermions in a random scalar potential, with $\mathsf{W}(p)\propto\text{\textbb{1}}$, the two curvatures coincide: then one just has
@@
We notice that in general $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}\neq-\mathsf{\Omega}^{\mathsf{W}}_{p_{\nu}p_{\mu}}$, at odds with the usual Berry curvature. Meanwhile, from Eq.(31) it appears that $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}$ is odd under TR and even under inversion, just like the usual Berry curvature.
Besides our rigorous formal proof of Eq.(27) (and of the energy current below) including semiclassical disorder corrections, equations Eqs. (31), (32) are the main new result of this work, displaying both qualitative and quantitative differences with the known theory for the case of fermions in a random scalar potential.
III.5 The disorder matrix structure $\hat{\sf W}$
We now provide more details on the interpretation and the determination in practice of the disorder matrix ${\mathsf{W}}(p)=\mathsf{\Gamma}\hat{W}(p)$, which plays a crucial role in defining the disorder-induced curvature $\mathsf{\Omega}^{\mathsf{W}}$ in Eqs.(32), (33).
An impurity localized at a given position modifies locally the hamiltonian. Should one place such an impurity at each site, translational invariance would be restored and the homogeneous hamiltonian would simply be shifted by a matrix function of momentum, $\mathsf{H}(p)\rightarrow\mathsf{H}(p)+\hat{W}(p)$. This defines $\hat{W}(p)$ and provides a way to compute it concretely. Such a disorder matrix can describe any local variations of the parameters of the hamiltonian (magnetic exchanges, hopping amplitudes, elastic moduli, etc), that define the matrix structure and momentum dependence of $\mathsf{H}(p)$: in fact, at this stage $\hat{W}(p)$ could simply be absorbed into a redefinition of $\mathsf{H}(p)$.
Now, more generally, in the inhomogeneous setting, distributing a “bunch” of impurities around position $X$ weighted by a smoothly decreasing function of distance shifts the hamiltonian as $\mathsf{H}(X,p)\rightarrow\mathsf{H}(X,p)+\tilde{W}(X,p)$, where $\tilde{W}(X,p)$ reduces to $\hat{W}(p)$ in the homogeneous limit where the impurities “extend” over the whole system. To describe the realistic case of correlated but dilute impurities, we describe them as a self-averaging, effectively random distribution. We introduce such disorder as a random variable, $\tilde{W}(X,p)\mapsto\hat{V}(X,p)$, and assume gaussian correlations between the impurity positions, Eq.(II.2), which equivalently in phase-space notations reads
$\langle\hat{V}_{ab}(X,k)\hat{V}_{cd}(X^{\prime},k^{\prime})\rangle=\hat{W}_{ab}(k)\hat{W}_{cd}(k^{\prime})\;w(X-X^{\prime}).$
Thus $\hat{V}(X,p)$ can be simply understood as a random (self-averaging) distribution of impurities with fixed matrix structure $\hat{W}(p)$, whose random positions $X$ have a gaussian correlation function $w(X-X^{\prime})$.
Although these formal details are needed to establish rigorously our results, we emphasize that determining $\hat{W}(p)$, whence ${\mathsf{W}}(p)=\mathsf{\Gamma}\hat{W}(p)$ and $\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}}$ from Eqs.(32), (33), in practice is more straightforward and can be done in a fully deterministic (non-random) setting.
@@ IV.1 Solving the kinetic equation
$\quad\times\delta\Big((\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p^{\prime})-(\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p)\Big),$ (36)
@@ IV.3 The “side-jump” accumulation term
Expanding the Moyal product, we find to the order $O(\hbar^{1})$ two new contributions in the energy current density. One of these is due to the perturbative shift of energy levels, which (see discussion below Eq.(III.1) and details in App. D.1) falls outside of the semiclassical picture: we do not consider it here. The other one can be identified as the side-jump current,
@@
$\qquad\qquad\qquad\times(p^{\prime}-p)_{\nu}\,(\mathsf{\Omega}^{\mathsf{W}}_{p_{\mu}p_{\nu}})_{n,n}(p)$
$\times\delta\Big((\underline{\mathsf{K}}_{\rm d})_{n,n}(p^{\prime})-(\underline{\mathsf{K}}_{\rm d})_{n,n}(p)\Big)\,(\underline{\mathsf{F}}_{\rm d})_{n,n}(X,p^{\prime}).$
@@
$\qquad\times\partial_{T}f((\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p))\;\partial_{p_{\lambda}}(\underline{\mathsf{K}}_{\rm d})_{n,n}(X,p).$
@@ IV.4 Total contribution of disorder to the thermal Hall conductivity
We note that because $\mathsf{\Omega}^{\mathsf{W}}$ is not antisymmetric a priori, ${J}_{x}^{\rm dis}\neq 0$ a priori, so that from Eqs. (37),(40) there may be a $O(\hbar^{1})$ contribution to the longitudinal thermal conductivity, which otherwise is mainly $O(\hbar^{0})$. In the limit of weak disorder, the $O(\hbar^{0})$ contribution is the dominant one, as it is larger by one power of the mean-free path.
IV.5 Analogous derivation for the fermionic Hall conductivities
Thermal Hall conductivity of fermions – The same steps can be adapted with almost no change to derive the intrinsic and disorder contributions to the thermal Hall effect of fermions (be they neutral or charged). Importantly, the normalization of eigenvectors is now $\mathsf{S}^{\dagger}\star\mathsf{S}=\text{\textbb{1}}$, so that $\underline{\mathsf{F}}_{\rm d}$ has dimension of a number not an energy. The equilibrium distribution is then $f(\varepsilon)=n_{\rm F}(\varepsilon)$, with $n_{\rm F}(x)=1/(e^{x/T}+1)$ the Fermi function. Then Eqs.(1), (2) still hold, up to the changes $n_{\rm B}\rightarrow n_{\rm F}$ and $\rm Tr_{+}\rightarrow Tr$, the trace over all fermionic bands.
Here it is worth noting, by inspection of the fermionic equivalents of Eqs.(1), (2), that $\kappa_{xy}^{\rm clean}$ gets contributions from all filled states (cf the $\Theta(\varepsilon-\underline{\mathsf{K}}_{d})$ factor), whereas $\kappa_{xy}^{\rm dis}$ gets contributions solely from states close to the Fermi level (cf the $\delta(\varepsilon-\underline{\mathsf{K}}_{d})$ factor). In particular, in a gapped fermionic phase, $\kappa_{xy}^{\rm dis}$ exhibits a thermal activation behavior and vanishes at temperatures well below the fermion gap, while no such behavior is expected for $\kappa_{xy}^{\rm clean}$. This makes it possible, in principle, to distinguish the two types of contributions from each other. We note that a similar separation of the contributions is not possible in the case of bosons.
As a direct application, it will be interesting to study thermal Hall transport e.g. in the quantum spin liquid phase of Kitaev’s honeycomb model with fermionic spinons in the presence of disorder.
Electrical Hall conductivity of fermions – The same steps can also be adapted with very little change to derive the electrical Hall effect of electrons. The equilibrium distribution is then $f(\varepsilon)=n_{\rm F}(\varepsilon-{\mathsf{e}}\phi)$, with $\mathsf{e}$ the electron charge and $\phi$ the electrical potential ($\bm{E}=-\bm{\nabla}\phi$ the electric field).
@@ IV.5 Analogous derivation for the fermionic Hall conductivities
$\partial_{t}{\rm Tr}\,\mathsf{F}=\frac{1}{\hbar}{\rm Im}{\rm Tr}\left[\mathsf{K}+\mathsf{V}\;\overset{\star}{,}\;\mathsf{F}\right],$ (42)
@@
$\times\left[{\mathsf{\Omega}}^{\mathsf{W}}_{p_{y}p_{\lambda}}(p)\,\partial_{p_{x}}\underline{\mathsf{K}}_{\rm d}(p)-{\mathsf{\Omega}}^{\mathsf{W}}_{p_{x}p_{\lambda}}(p)\,\partial_{p_{y}}\underline{\mathsf{K}}_{\rm d}(p)\right],$ (43)
see details in Appendix F.2. This formula generalizes the known result for electron side-jump based on a wavepacket analysis Sinitsyn et al. (2005), by allowing for impurities with a local matrix structure $\hat{W}$. In the limit case of fermions in a random scalar potential (where $\mathsf{\Omega}^{\mathsf{W}}=\mathsf{\Omega}$ reduces to the Berry curvature), assuming a quadratic dispersion and a uniform (i.e. momentum-independent) Berry curvature $\Omega^{z}$, one recovers immediately the familiar result $\sigma_{xy}^{\rm dis}=-2\sigma_{xy}^{\rm clean}$, where $\sigma_{xy}^{\rm clean}={\mathsf{e}}^{2}p_{\rm F}^{2}\Omega^{z}/4\pi$ is given by the Karplus-Luttinger formula ($p_{\rm F}$ the Fermi momentum).
IV.6 Generalizations and discussion
The above calculations can be generalized to the case of several independent sets of impurities, labeled $(\iota)$, each with its own matrix structure $\hat{W}^{(\iota)}$ and spatial correlation function $w^{(\iota)}$. One can also include the effect of interactions as an extra collision integral,
$\left(\text{d}_{t}\,\underline{\mathsf{F}}_{\rm d}\right)_{\rm coll}^{\rm int}(X,p)=\int_{p^{\prime}}\Upsilon^{\rm int}_{p,p^{\prime}}\left(\mathsf{g}_{\rm d}(X,p^{\prime})-\mathsf{g}_{\rm d}(X,p)\right),$ (44)
to be added on the right-hand side of Eq.(III.3), and assumed to be diagonal in the (implicit) band index $n$.
Details of this generalization are provided in Appendix E.4. The upshot is that our main result, Eq.(2), still holds up to the replacement
[…]
By allowing for complex realistic systems with various types of impurities and a minimal inclusion of many-body interaction effects, Eq.(45) plugged into Eq.(2) accounts for many pieces of physics that are absent from the clean contribution Eq.(1).
That the contribution we derived in Eq.(2) is non-perturbative in the disorder strength (i.e. it scales as $n_{\rm imp}^{0}$) is an intriguing, yet well-known Nagaosa et al. property of coordinate shift effects in anomalous Hall transport, and may be given a very intuitive explanation. Whenever heat is carried by well-defined quasiparticles, one may write $\kappa\sim c\,v\,l$ with $c$ the specific heat, $v$ the velocity and $l$ the mean-free path. This holds for any contribution independently, so focussing on the side-jump Hall conductivity,
$\kappa_{xy}^{\rm sj}\sim c\,v_{\perp}^{\rm sj}\,l,$ (46)
where $v_{\perp}^{\rm sj}$ is now the side-jump transverse velocity, $v_{\perp}^{\rm sj}\propto n_{\rm imp}$. In Eq.(2) we have assumed that $l$ is governed by disorder scattering, $l^{-1}\propto n_{\rm imp}$, which entails that $\kappa_{xy}^{\rm sj}$ is independent of $n_{\rm imp}$. More generally, when interactions also contribute to the mean-free path, $\kappa_{xy}^{\rm sj}\propto l_{\rm net}/l_{\rm imp}$: the ratio of the net mean-free path (including all effects) to the impurity mean-free path. Precisely this correction is captured in Eq.(45).
We note that a recent effort was made to include the effect of band broadening in the intrinsic thermal Hall effect Koyama and Nasu : this effect is perturbative in the disorder strength, and thus fully distinct from side-jump which appears as a separate, non-perturbative contribution.
@@ V Applications
We first provide a reminder of the equivalence between our formalism, introduced in Sec.II.1, and general bosonic lattice models expressed in a momentum-space formulation. As a direct concrete application, we consider the honeycomb-lattice $K\Gamma\Gamma^{\prime}$ model in a field, in the presence of disorder. Then, for a different illustration of our results and a more formal discussion, we adopt a more “low-energy” perspective, and consider an application to three slightly different physical models with disorder, the results of which are displayed in Fig. 1.
V.1 General bosonic hamiltonians in second-quantized notations
A homogeneous quadratic theory of bosons, in its most general form in momentum space, reads
$H_{\rm hom}=\frac{\hbar}{2}\int_{\bm{k}}\left({\sf b}^{\dagger}_{i,\bm{k}}\,,\,{\sf b}_{i,-\bm{k}}\right)\begin{pmatrix}(A_{\bm{k}})_{ij}&(B_{\bm{k}})_{ij}\\ (B_{-\bm{k}})^{*}_{ij}&(A^{\top}_{-\bm{k}})_{ij}\end{pmatrix}\begin{pmatrix}{\sf b}_{j,\bm{k}}\\ {\sf b}^{\dagger}_{j,-\bm{k}}\end{pmatrix}$ (47)
where ${\sf b}^{\dagger}_{i,\bm{k}}$ is a creation operator for a complex boson with flavor index $i$ and momentum $\bm{k}$. Hermiticity imposes $A_{\bm{k}}=A^{\dagger}_{\bm{k}}$ and without loss of generality $B_{\bm{k}}=B^{\top}_{-\bm{k}}$. Real bosons may then be defined as
$\Phi^{\top}_{i,\bm{k}}=\tfrac{1}{\sqrt{2\hbar}}\left({\sf b}^{\dagger}_{i,\bm{k}}+{\sf b}_{i,-\bm{k}}\,,\,i({\sf b}^{\dagger}_{i,\bm{k}}-{\sf b}_{i,-\bm{k}})\right),$ (48)
with two components $\gamma=1,2$ commuting as $\left[\Phi^{1}_{i,\bm{k}},\Phi^{2}_{j,\bm{k}^{\prime}}\right]=i\hbar\,\delta_{i,j}\delta_{\bm{k}+\bm{k}^{\prime}}$. Defining the joint index $a=(\gamma,i)$ so that $\Phi^{\top}_{a,\bm{k}}=\left(\{\Phi^{1}_{i,\bm{k}}\}_{i},\{\Phi^{2}_{i,\bm{k}}\}_{i}\right)_{a=(\gamma,i)}$, upon Fourier transformation $H_{\rm hom}$ may be rewritten precisely as Eq.(3) with
$\mathsf{H}_{ab}(x_{1},x_{2})$ $=\int_{\bm{k}}e^{-i\bm{k}(\bm{x}_{1}-\bm{x}_{2})}\begin{pmatrix}C^{+}_{\bm{k}}&D^{+}_{\bm{k}}\\ D^{-}_{\bm{k}}&C^{-}_{\bm{k}}\end{pmatrix}_{ab}$ (49)
and
$C^{\pm}_{\bm{k}}$ $=\tfrac{1}{2}(A_{\bm{k}}+A^{*}_{-\bm{k}})\pm\tfrac{1}{2}(B_{\bm{k}}+B^{*}_{-\bm{k}}),$ (50a)
$D^{\pm}_{\bm{k}}$ $=-\tfrac{i}{2}(B_{\bm{k}}-B^{*}_{-\bm{k}})\pm\tfrac{i}{2}(A_{\bm{k}}-A^{*}_{-\bm{k}}).$ (50b)
From $(C^{\pm}_{\bm{k}})^{\top}=C^{\pm}_{-\bm{k}}$ and $(D^{\pm}_{\bm{k}})^{\top}=D^{\mp}_{-\bm{k}}$ one can check ${\sf H}_{ab}(x_{1},x_{2})={\sf H}_{ba}(x_{2},x_{1})$. From $(C^{\pm}_{\bm{k}})^{*}=C^{\pm}_{-\bm{k}}$ and $(D^{\pm}_{\bm{k}})^{*}=D^{\pm}_{-\bm{k}}$ one can check ${\sf H}_{ab}(x_{1},x_{2})\in\mathbb{R}$. Besides, with this choice of basis, the commutation relation between fields $\Phi_{a}(x)$ is just given by
${\sf\Gamma}(x_{1},x_{2})=\begin{pmatrix}0&i\,\text{\textbb{1}}\\ -i\,\text{\textbb{1}}&0\end{pmatrix}\,\delta(x_{1}-x_{2}).$ (51)
Thus any quadratic theory written in terms of ${\sf b}^{\dagger}_{i,\bm{k}},{\sf b}_{i,\bm{k}}$ can equivalently be put into the form introduced in Sec.II.1. This includes, notably, any spin hamiltonian defined on any lattice and expanded into magnons to the quadratic order.  [^5] Note that here the theory is homogeneous (i.e. $\mathsf{H}$ only depends on the relative coordinate $x_{1}-x_{2}$), but it can be made inhomogeneous a posteriori by allowing the coefficients ${\sf H}_{ab}$ to depend on $(x_{1}+x_{2})/{2}$ and reintroducing the Moyal product.
In the case of canonical bosons, such as phonons or magnons in the spin path integral representation Auerbach (1994), the Berry curvature is known (see Qin et al. and discussion of the $d_{i}$ coefficients in Eq.(V.3.1)) to arise from diagonal elements of the dynamical matrix $\mathsf{K}=\mathsf{\Gamma}\mathsf{H}$. These are proportional to the $D^{\pm}_{\bm{k}}$ matrix elements: in models where $D^{\pm}_{\bm{k}}=0$, no intrinsic THE is possible. Now, for such canonical bosons, the TR operation is $\mathsf{b}_{\bm{k}}{}^{(}{}^{\dagger}{}^{)}\rightarrow\mathsf{b}_{-\bm{k}}{}^{(}{}^{\dagger}{}^{)}$: this implies the symmetries $A_{\bm{k}}\overset{\rm TR}{=}A^{*}_{-\bm{k}}$ and $B_{\bm{k}}\overset{\rm TR}{=}B^{*}_{-\bm{k}}$. Thus $D^{\pm}_{\bm{k}}$ involves precisely those combinations of $A_{\bm{k}},B_{\bm{k}}$ that break TR.  [^6] The situation is less transparent for non-canonical bosons, for instance the Holstein-Primakoff representation where TR acts as $\mathsf{b}_{\bm{k}}\rightarrow-\mathsf{b}_{-\bm{k}}^{\dagger},\mathsf{b}_{\bm{k}}^{\dagger}\rightarrow-\mathsf{b}_{-\bm{k}}$. This illustrates the general fact that TR breaking is a necessary condition for a THE.
The same procedure and physical arguments hold for the disorder matrix $\hat{W}(p)$. Namely, the homogeneous disorder hamiltonian as described in Sec.III.5 can be written most generally as
$H_{\rm dis}=\frac{\hbar}{2}\int_{\bm{k}}\left({\sf b}^{\dagger}_{i,\bm{k}}\,,\,{\sf b}_{i,-\bm{k}}\right)\begin{pmatrix}(\tilde{A}_{\bm{k}})_{ij}&(\tilde{B}_{\bm{k}})_{ij}\\ (\tilde{B}_{-\bm{k}})^{*}_{ij}&(\tilde{A}^{\top}_{-\bm{k}})_{ij}\end{pmatrix}\begin{pmatrix}{\sf b}_{j,\bm{k}}\\ {\sf b}^{\dagger}_{j,-\bm{k}}\end{pmatrix}$ (52)
where $\tilde{A}_{\bm{k}}=\tilde{A}^{\dagger}_{\bm{k}}$ and $\tilde{B}_{\bm{k}}=\tilde{B}^{\top}_{-\bm{k}}$ can be chosen independently from $A_{\bm{k}}$ and $B_{\bm{k}}$. Then, in these notations
$\hat{W}(k)=\begin{pmatrix}\tilde{C}^{+}_{\bm{k}}&\tilde{D}^{+}_{\bm{k}}\\ \tilde{D}^{-}_{\bm{k}}&\tilde{C}^{-}_{\bm{k}}\end{pmatrix}$ (53)
where $\tilde{C}^{\pm}_{\bm{k}},\tilde{D}^{\pm}_{\bm{k}}$ are defined in analogy to Eqs.(50a),(50b). This determines $\mathsf{W}=\mathsf{\Gamma}\hat{W}$.
Once $H_{\rm hom}$ and $H_{\rm dis}$ are specified, it is straightforward to obtain both $\mathsf{\Omega}_{p_{x}p_{y}}$ and $\tfrac{1}{2}(\mathsf{\Omega}^{\mathsf{W}}_{p_{x}p_{y}}-\mathsf{\Omega}^{\mathsf{W}}_{p_{y}p_{x}})$: we remind the reader that a recipe suitable for numerical evaluation is detailed in App.G.
V.2 Concrete application: honeycomb $K\Gamma\Gamma^{\prime}$ model in a magnetic field with disorder
We now appy the above to a particular case of a bosonic Hamiltonian, the ferromagnetic honeycomb-lattice $K\Gamma\Gamma^{\prime}$ spin model in a field, whose dynamics in the field-aligned ordered phase can be described in terms of magnon excitations. The clean thermal Hall conductivity of magnons in this model was extensively studied in Ref. Chern et al. (2021), which discusses in great detail how the interplay of globally broken TR symmetry and anisotropic magnetic exchange generates a magnon Berry curvature and a sizable THE. We will not reproduce this discussion here, but investigate the role of disorder as random local variations of the parameters of the model.
The microscopic spin hamiltonian reads
$H_{\rm spin}=\sum_{\gamma\in\{x,y,z\}}\sum_{\langle\mathsf{i},\mathsf{j}\rangle\in\gamma}\mathbf{S}_{\mathsf{i}}^{\top}H_{\gamma}\,\mathbf{S}_{\mathsf{j}}-\sum_{\mathsf{i}}\bm{h}\cdot\mathbf{S}_{\mathsf{i}},$ (54)
where $\langle\mathsf{i},\mathsf{j}\rangle\in\gamma$ labels neighboring sites $\mathsf{i},\mathsf{j}$ sharing a bond of type $\gamma\in\{x,y,z\}$, the field $\bm{h}$ points along the $\hat{a}$ axis (for geometrical conventions see Fig. 1 in Ref. Chern et al. (2021)), $\mathbf{S}_{\mathsf{i}}$ is the spin operator at site $\mathsf{i}$, and $H_{x},H_{y},H_{z}$ are $3\times 3$ matrices with elements
$(H_{\gamma})_{\mu\nu}$ $=J\,\delta_{\mu\nu}+K\,\delta_{\mu\gamma}\delta_{\nu\gamma}+\Gamma\,|\epsilon_{\mu\nu\gamma}|$
$\quad+\Gamma^{\prime}\,(\delta_{\mu\gamma}+\delta_{\nu\gamma})(1-\delta_{\mu\gamma}\delta_{\nu\gamma})$ (55)
indexed by $\mu,\nu\in\{x,y,z\}$. As in Ref. Chern et al. (2021), we consider the pure $K\Gamma\Gamma^{\prime}$ model where $J=0$, in the ferromagnetic case where $K<0$. Upon spin wave expansion of the $\mathbf{S}_{\mathsf{i}}$ operators around the classical field-polarized state, $H_{\rm spin}$ can be recast into the form of Eq.(47) where the $2\times 2$ matrices $A_{\bm{k}},B_{\bm{k}}$ are provided in Ref. Chern et al. (2021), equations (S3b–c) and (S7a) – again we do not reproduce the details here and refer the reader to the original paper.
For real materials, all parameters in the model depend on the chemical environment of the spins $\mathbf{S}_{\mathsf{i}}$: this holds for magnetic exchange parameters $J,K,\Gamma,\Gamma^{\prime}$ as well as for $\bm{h}={\sf g}\mu_{\rm B}\bm{B}$ through the g-tensor $\sf g$ (here $\bm{B}$ is the applied magnetic field). Since this chemical environment may be affected by disorder, all parameters $(J,K,\Gamma,\Gamma^{\prime},{\sf g})$ have a random component, which can be described by a disorder Hamiltonian in the form of Eq.(52). In the following we will consider these various instances of disorder independently from each other, and call “$X$-type disorder” (where $X\in\{J,K,\Gamma,\Gamma^{\prime},{\sf g}\}$) that instance of $H_{\rm dis}$ where $\tilde{A}_{\bm{k}},\tilde{B}_{\bm{k}}$ are derived from $A_{\bm{k}},B_{\bm{k}}$ by setting all parameters to zero except $X\rightarrow 1$. We note that because the computed side-jump effect does not depend on the disorder strength, any global rescaling of $H_{\rm dis}$ leads to the same results.
Using the procedure detailed in App.G, we evaluated the thermal Hall conductivity given by Eqs.(1) ($\kappa_{xy}^{\rm clean}$, intrinsic contribution) and (2) ($\kappa_{xy}^{\rm dis}$, disorder contribution) for $K$-type, $\Gamma^{\prime}$-type  [^7] Because $\Gamma$-type disorder does not satisfy the requirement stated below Eq.(33), it cannot be studied within our small-deviation expansion., $J$-type and $\sf g$-type disorder. We used the same relative parameter values $(K,\Gamma,\Gamma^{\prime})=(-1,0.2,-0.02)$ and $h/S|K|=0.1$ as in Ref. Chern et al. (2021) (with $S=1/2$). We also used the same numerical values in physical units, namely $|K|\approx 80~\rm K$, $\sf g\approx 2.3$ hence $B\approx 3~\rm T$, and interlayer spacing $\kappa_{xy}^{2D}/\kappa_{xy}^{3D}=5.72\,\AA$ , all inspired by $\alpha$-RuCl3. The results are displayed in Fig. 3.
Figure 3: Thermal Hall conductivity of magnons in the disordered $K\Gamma\Gamma^{\prime}$ model. Blue, red and yellow curves represent the intrinsic, extrinsic and total contribution, respectively. Different symbols account for different instances of disorder (for instance filled disks mean $K$-type disorder). Right axis: experimental quantity $\kappa_{xy}^{3D}/T$ in SI units. Left axis: equivalent 2D quantity $\kappa_{xy}^{2D}/T$ in units of the thermal Hall quantum.
We find that different types of disorder can produce very different extrinsic contributions to the THE, that are typically of the same order of magnitude as the intrinsic contribution, and that can be of the same or the opposite sign (depending in particular on temperature). In turn, the total thermal Hall conductivity $\kappa_{xy}^{\rm tot}=\kappa_{xy}^{\rm clean}+\kappa_{xy}^{\rm dis}$ is generally quite different from the sole intrinsic contribution, confirming our qualitative discussions in previous sections. Qualitative differences appear: notably, we find that $\kappa_{xy}^{\rm tot}$ can undergo a sign change as a function of $T$ that is absent in $\kappa_{xy}^{\rm clean}$. There are also clear quantitative differences: in some temperature ranges, $\kappa_{xy}$ can be enhanced by a $O(1)$ factor, that for some types of disorder (especially $\sf g$-type) can be as large as 2 or more.
We note that for the latter type of disorder, our quantitative estimates for $\kappa_{xy}/T$ are comparable to those found in experiments on $\alpha$-RuCl3 Czajka et al. (2023), although for different experimental parameters $(B,T)$, and are interestingly close to the half-quantized value that was predicted Vinkler-Aviv and Rosch (2018); Ye et al. (2018) to arise from very different excitations. We do not claim, from our simple study, any quantitative description of that compound, which was argued to be rather described by a $JK\Gamma\Gamma^{\prime}$ Kim and Kee (2016) or a $JK\Gamma J_{3}$ Winter et al. (2016); Wang et al. (2017); Winter et al. (2017) model. However, our results clearly show that side-jump effects produce an indispensable contribution to the total thermal Hall effect, that should always be taken into account in order to reach quantitative agreement with the experimental data. To achieve this, a realistic description of disorder in the actual compound, that may involve different types of impurities (that can be incorporated as described in Sec.IV.6), will also be necessary.
V.3 Formal application and discussion in a low-energy bosonic model
V.3.1 Introduction and realization in physical systems